Why is theoretical probability called theoretical probability?
It is the theory of what might happen, but not actually what happens. In theory, if you spin a coin 100 times, it should come up on heads 50 times, as there is a 1 in 2 chance of you getting heads on each spin. If you actually do spin a coin 100 times, the total of heads is the experimental probability, so what you actually get. That may not be 50. It is likely to be close to 50 though.
What is the probability of rolling a 3?
assuming a single six sided die the probability of rolling a three is 1/6
Why mean and standard deviation are used for inferential statistics?
The following are the two main reasons.
I am not sure but I think 3/36
The difference between the theoretical and empirical probability?
The term empirical means "based on observation or experiment." An empirical probability is generally, but not always, given with a number indicating the possible percent error (e.g. 80+/-3%). A theoretical probability, however, is one that is calculatedbased on theory, i.e., without running any experiments.
Since there is no theory that will calculate the probability that an area will experience an earthquake within a given time frame, the 90% figure is an empirical probability, presumably based on data of major earthquakes in the San Francisco area over past years.
2 examples how the theory of probability is applied in real life situations?
how theory of probability used in real life
What is a probability of drawing a club and a 7?
drawing a club
total club =13
P(Club)= 13/52
drawing a 7
total 7 = 4
P(7) = 4/52
P(a club and a 7) = P(Club) + P(7) as events are mutualy exculisve
P(a club and a 7) = 13/52+4/52
P(a club and a 7) = 17/52
If you roll a fair die what is the theoretical probability of getting a 5?
Since there are 6 numbers on a die (1-6), then the probability of rolling a 5 would be 1 out of 6.
What is the probability of having a girl versus having a boy?
The probability of having a girl versus a boy is 1/2 because there is two things you have a chance of getting and you can only get one or the other.
You roll a die twice what is the probability of rolling a 2 then a 3?
Prob(2 and then a 3) = Prob(2 on roll 1)*Prob(3 on roll 2) since the events are independent
= 1/6 * 1/6 = 1/36 = 0.0277... or 2.77 %
Find the number of distinguishable permutations of the letters in the word Cincinnati?
There are ten letters in the word. The total number of possible permutations is
(10) x (9) x (8) x (7) x (6) x (5) x (4) x (3) x (2) = 3,628,800
But the two 'c's can be arranged in either of 2 ways with no distinguishable change.
Also, the three 'i's can be arranged in any of (3 x 2) = 6 ways with no distinguishable change.
And the three 't's can be arranged in any of (3 x 2) = 6 ways with no distinguishable change.
So the total number of possible permutations can be divided by (2 x 6 x 6) = 72, the number of
times each distinguishable permutation occurs with different and indisnguishable arrangements
of 'c', 'i', and 't'.
We're left with
(10) x (9) x (8) x (7) x (...) x (5) x (...) x (...) x (2) = (3,628,800/72) = 50,400 distinguishable arrangements.
What is the probability of getting a head if coin is tossed 5 times?
The probability of getting a head on each occasion if a coin is tossed five times is equal to 1/2 x 1/2 x 1/2 x 1/2 x 1/2 (that is, 1/25) = 1/32 or one in thirty-two.
The probability of getting a head on the fifth toss is 1/2. All other tosses are independent of the fifth toss.
The probability of getting at least one head when tossing the coin 5 times: Since there are 32 possible combinations and only one of those combinations is all tails you could expect to toss a combination with a head 31/32 times. The probability is 0.96875.
95%
five minutes
What is the probability of getting 5 or 6 on a single throw?
There are 6 outcomes on a die roll (1,2,3,4,5,6); 5 & 6 constitutes 2 of them so the probability is 2/6 or 1/3.
The probability is 7,893,600/311,875,200 = 0.0253
a large number of samples of size 50 were selected at random from a normal population with mean and variance.The mean and standard error of the sampling distribution of the sample mean were obtain 2500 and 4 respectivly.Find the mean and varince of the population?
No, it is not true. Probability can be used to describe some properties of the variation but not all.
How many different 3 letter permutations can be formed from the letters in the word count?
There are 5*4*3 = 60 permutations.